# On a certain subclass of strongly starlike functions

### v3

@@ Abstract.
$-\frac{\pi\alpha_{1}}{2}<\arg\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}<\frac{\pi\alpha_{2}}{2},\quad(z\in\mathbb{D}).$
@@ 1. Introduction
Strongly starlike functions are a subset of starlike functions, where the condition is stronger, ensuring a more restricted behavior of the image of the unit disk under the function. These functions are of interest in complex analysis and have applications in various areas, including the study of univalent functions and geometric function theory.
@@ 2. Preliminaries
Consider the set $\mathcal{H}$ consisting of functions, denoted $f$, that are holomorphic within the open unit disk $\mathbb{D}$, defined as $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$. Furthermore, let $\mathcal{A}$ be a subset of $\mathcal{H}$, containing functions, also denoted as $f$, which can be represented in the form:
@@
These functions in $\mathcal{A}$ are subject to a normalization condition, namely, $f(0)=0$ and $f^{\prime}(0)=1$, when evaluated within the unit disk $\mathbb{D}$. The subclass of $\mathcal{A}$ consisting of all univalent functions in $\mathbb{D}$ will be denoted by $\mathcal{S}$. We say that a function $f\in\mathcal{S}$ is starlike if, and only if,
@@
In the open unit disk $\mathbb{D}$, two analytic functions, denoted as $f$ and $g$, are considered in a subordinate relationship if $f(z)\prec g(z)$ or simply $f\prec g$. This relationship holds true when there exists an analytic function $w$ defined within $\mathbb{D}$, satisfying the initial conditions $w(0)=0$ and $|w(z)|\leq 1$, and such that $f(z)=g(w(z))$ for all $z\in\mathbb{D}$.
In this paper, we consider the following analytic function $G(z):=G(\alpha_{1},\alpha_{2},c)(z)$:
(2.2) $G(\alpha_{1},\alpha_{2},c)(z):=\left(\frac{1+cz}{1-z}\right)^{(\alpha_{1}+\alpha_{2})/2},\quad(z\in\mathbb{D}),$
@@
$\left\{w\in\mathbb{C}:-\frac{\pi\alpha_{1}}{2}<\arg\,w<-\frac{\pi\alpha_{2}}{2}\right\}.$
See [35, Lemma 2] for more details. It is also worth mentioning that this function is a special case of the function defined in [1, (2.2)], where $x=c$ and $\alpha=(\alpha_{1}+\alpha_{2})/2$. It follows from [1, (2.2)] that
$G(z)=1+\sum_{n=1}^{\infty}A_{n}(\alpha_{1},\alpha_{2},c)z^{n},$
@@
(2.3) $A_{n}:=A_{n}(\alpha_{1},\alpha_{2},c)=\sum_{k=1}^{n}\binom{n-1}{k-1}\binom{(\alpha_{1}+\alpha_{2})/2}{k}(1+c)^{k}=\frac{\alpha_{1}+\alpha_{2}}{2}(1+c)B_{n}(\alpha_{1},\alpha_{2},c),$
with
(2.4) $B_{n}(\alpha_{1},\alpha_{2},c):=F(1-n,1-(\alpha_{1}+\alpha_{2})/2;2;1+c),\quad(n=1,2,\ldots).$
@@
$F(a,b;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{(c)_{n}(1)_{n}}z^{n},\quad(a,b,c\in\mathbb{C},c\not\in\{0,-1,-2,\ldots\},z\in\mathbb{D}).$
The expression $(a)_{n}$, denoted as the Pochhammer symbol, signifies the product of consecutive terms starting from $a$ and ending at $a+n-1$, i.e., $(a)_{n}=a(a+1)\cdots(a+n-1)$ with the convention that $(a)_{0}$ is equal to $1$. For a deeper exploration of $G$ and its properties, refer to the comprehensive discussion in reference [1, Section 2].
@@
$|A_{n}|\leq(\alpha_{1}+\alpha_{2})\cos(\pi\theta/2)|B_{n}(\alpha_{1},\alpha_{2},1)|,$
where $\theta=(\alpha_{2}-\alpha_{1})/(\alpha_{2}+\alpha_{1})$.
The main purpose of this paper is to study the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ which is provided below:
@@
(2.5) $-\frac{\pi\alpha_{1}}{2}<\arg\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}<\frac{\pi\alpha_{2}}{2},\quad(z\in\mathbb{D}),$
@@
The introduction of the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ is attributed to Takahashi and Nunokawa, specifically in the context of inequalities (2.5), see [33]. We recall here the fact, that in [7] and in [8] a similar class was studied. It is clear that $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})\subset\mathcal{S}^{*}$ and that $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ is a subclass of the class of strongly starlike functions of order $\beta=\max\{\alpha_{1},\alpha_{2}\}$, i.e. $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})\subset\mathcal{S}^{*}(\beta,\beta)\equiv\mathcal{SS}^{*}(\beta)$.
@@
Following is another useful lemma (see [25]).
Lemma 2.4.
@@
$|w_{2}-tw_{1}^{2}|\leq\left\{\begin{array}[]{ll}-t,&\hbox{$t\leq-1$;}\\ \\ 1&\hbox{$-1\leq t\leq 1$;}\\ \\ t,&\hbox{$t\geq 1$.}\end{array}\right.$
All inequalities are sharp.
@@
(See [19]) If $w(z)=\sum_{k=1}^{\infty}w_{k}z^{k}$ is a Schwarz function, then for any complex numbers $\rho$ and $\tau$ the following sharp estimate holds:
@@
(2.8) $H(\rho,\tau)=\left\{\begin{array}[]{ll}1&\hbox{for $(\rho,\tau)\in\Omega_{1}\cup\Omega_{2}$}\\ \\ |\rho|&\hbox{for $(\rho,\tau)\in\bigcup_{k=3}^{7}\Omega_{k}$;}\\ \\ \frac{2}{3}(|\rho|+1)\left(\frac{|\rho|+1}{3(|\rho|+1+\tau)}\right)^{\frac{1}{2}}&\hbox{for $(\rho,\tau)\in\Omega_{8}\cup\Omega_{9}$}\\ \\ \frac{\tau}{3}\left(\frac{\rho^{2}-4}{\rho^{2}-4\tau}\right)^{\frac{1}{2}}&\hbox{for $(\rho,\tau)\in\Omega_{10}\cup\Omega_{11}\setminus\{\pm 2,1\}$}\\ \\ \frac{2}{3}(|\rho|-1)\left(\frac{|\rho|-1}{3(|\rho|-1-\tau)}\right)^{\frac{1}{2}}&\hbox{for $(\rho,\tau)\in\Omega_{12}.$}\\ \\ \end{array}\right.$
@@ 3. Main Results
In this paper, we find a general upper bound for $|a_{n}|$, where $n=2,3,\ldots$. Our findings generally do not produce sharp results, except when $n=2$. But in Theorem 3.4, we obtain sharp estimates for the first three coefficients.
@@
Let $f$ be of the form (2.1) belonging to the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$. Then
(3.1) $|a_{n}|\leq\left\{\begin{array}[]{ll}(\alpha_{1}+\alpha_{2})\cos(\pi\theta/2),&\qquad\hbox{$n=2$;}\\ \\ \frac{\alpha_{1}+\alpha_{2}}{n-1}\cos(\pi\theta/2)\prod_{k=2}^{n-1}\left(1+\frac{\alpha_{1}+\alpha_{2}}{k-1}\cos(\pi\theta/2)\right),&\qquad\hbox{$n=3,4,\ldots$,}\end{array}\right.$
where $c=e^{\pi i\theta}$, $\theta=(\alpha_{2}-\alpha_{1})/(\alpha_{2}+\alpha_{1})$ and $0<\alpha_{1},\alpha_{2}\leq 1$. The inequality is sharp only for $n=2$.
@@
(3.4) $|Q_{n}|\leq|A_{1}|=(\alpha_{1}+\alpha_{2})\cos(\pi\theta/2),\quad(n=1,2,3,\ldots).$
@@
$\leq\frac{\alpha_{1}+\alpha_{2}}{n-1}\cos(\pi\theta/2)(|a_{1}|+|a_{2}|+\cdots+|a_{n-1}|)$
$=\frac{\alpha_{1}+\alpha_{2}}{n-1}\cos(\pi\theta/2)\sum_{k=1}^{n-1}|a_{k}|.$
It is clear that $|a_{2}|\leq(\alpha_{1}+\alpha_{2})\cos(\pi\theta/2)$. To prove the remaining part of the theorem, we need to show that
(3.5) $\frac{\alpha_{1}+\alpha_{2}}{n-1}\cos(\pi\theta/2)\sum_{k=1}^{n-1}|a_{k}|\leq\frac{\alpha_{1}+\alpha_{2}}{n-1}\cos(\pi\theta/2)\prod_{k=2}^{n-1}\left(1+\frac{\alpha_{1}+\alpha_{2}}{k-1}\cos(\pi\theta/2)\right),$
for $n=3,4,\ldots$. Using induction and simple calculation, we could prove the inequality (3.5). Hence, the desired estimate for $|a_{n}|$, $(n=3,4,5,\ldots)$, follows as asserted in (3.1). This completes the proof of the theorem. ∎
@@
where $0<\beta\leq 1$. The first inequality is sharp.
@@
Consider a function $f\in\mathcal{A}$, which takes the form described by equation (2.1) and belongs to the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$, where $\alpha_{1},\alpha_{2}\in(0,1]$. Then
$|a_{2}|\leq(\alpha_{1}+\alpha_{2})\cos\frac{\pi}{2}\theta,\qquad|a_{3}|\leq\left(1+(\alpha_{1}+\alpha_{2}-1)(1+e^{i\pi\theta})\right)\frac{(\alpha_{1}+\alpha_{2})}{2}\cos\frac{\pi}{2}\theta,$
@@
$|a_{4}|\leq\frac{\alpha_{1}+\alpha_{2}}{3}\,H(\rho,\tau),$
where $H(\rho,\tau)$ is given by (2.8). All inequalities are sharp.
@@
where $A_{n}$, $n=1,2,3$, defined in (2.3). Equating the coefficients of the corresponding terms in the last relation leads to
@@
It follows from the first equality of (3.7) that $|a_{2}|\leq(\alpha_{1}+\alpha_{2})\cos(\pi\theta/2)$. From the second equality of (3.7) and Lemma 2.5, we get
$2|a_{3}|\leq|A_{1}|\left|w_{2}-\left(-\frac{A_{2}}{A_{1}}-A_{1}\right)w_{1}^{2}\right|\leq|A_{1}|\left(\frac{A_{2}}{A_{1}}+A_{1}\right).$
It is easy to see that
$\frac{A_{2}}{A_{1}}+A_{1}=1+(\alpha_{1}+\alpha_{2}-1)(1+e^{i\pi\theta}),\quad\theta=\frac{\alpha_{2}-\alpha_{1}}{\alpha_{1}+\alpha_{2}},$
which gives the bound for $|a_{3}|$. From the coefficient relation (3.8)
@@
$\rho=\tfrac{3}{2}A_{1}+\frac{2A_{2}}{A_{1}},\quad\tau=\tfrac{1}{2}A_{1}^{2}+\tfrac{3}{2}A_{2}+\frac{A_{3}}{A_{1}},$
@@
$|a_{4}|\leq\frac{A_{1}}{3}\,H(\rho,\tau)=\frac{\alpha_{1}+\alpha_{2}}{3}\,H(\rho,\tau),$
where $H(\rho,\tau)$ is given by (2.8). ∎
@@
We summarize the bounds of $|a_{4}|$ for some special choices of $\alpha_{1},\alpha_{2}$:
Case 1. For $\alpha_{1}=\alpha_{2}=1$, we have $A_{1}=2$, $A_{2}=2.5$, $A_{3}=3.5$, and
$\rho=5,\ \tau=6,5.$
Since $(\rho,\tau)\in\Omega_{7}$, $H(\rho,\tau)=|\rho|=5$. Thus,
$|a_{4}|\leq\frac{2}{3}\cdot 5=\frac{10}{3}.$
Case 2. For $\alpha_{1}=\alpha_{2}=1/2$, we have $A_{1}=1$, $A_{2}=0.5$, $A_{3}=0.375$, and
$\rho=2.5,\ \tau=1.625.$
Since $(\rho,\tau)\in\Omega_{6}$, $H(\rho,\tau)=|\rho|=2.5$. Thus,
$|a_{4}|\leq\frac{1}{3}\cdot 2.5=\frac{5}{6}\approx 0.833.$
Case 3. For $\alpha_{1}=1,\alpha_{2}=0$ or $\alpha_{1}=0,\alpha_{2}=1$, we have $A_{1}=1$, $A_{2}=0.25$, $A_{3}=0.125$, and
$\rho=2,\ \tau=1.$
Since $(\rho,\tau)\in\Omega_{2}$, $H(\rho,\tau)=1$. Thus,
$|a_{4}|\leq\frac{1}{3}\cdot 1=\frac{1}{3}\approx 0.333.$
Case 4. For $(\alpha_{1},\alpha_{2})\to(0,0)$, we have $\rho,\tau\to 0$ and $(\rho,\tau)\in\Omega_{1}$, $H(\rho,\tau)=1$. Thus,
$|a_{4}|\leq\frac{1}{3}\cdot 1=\frac{1}{3}\approx 0.333.$
These bounds are sharp, attained by the extremal functions described in Lemma 2.6.
@@
$|\gamma_{1}|\leq 1\quad{\rm and}\quad|\gamma_{2}|\leq\frac{1}{2}(1+2e^{-2})\approx 0.635\ldots,$
hold. For $n\geq 3$, the estimate of $\gamma_{n}$ is much harder and no significant upper bounds for $|\gamma_{n}|$ were obtained when $f\in\mathcal{S}$ and is still open for $n\geq 3$. The sharp upper bounds for the modulus of logarithmic coefficients have been established for only a limited number of function subclasses within $\mathcal{S}$, the class $\mathcal{S}^{*}$ offers a relatively straightforward proof that $|\gamma_{n}|\leq 1/n$ for $n\geq 1$, with equality observed in the Koebe function. For other subclasses within $\mathcal{S}$, further details can be found in references [10, 20, 21, 34].
In the subsequent discussion, we turn our attention to the task of estimating the logarithmic coefficients for functions belonging to the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$. First, we need the following theorem:
@@
where $G(z)-1$ is convex univalent. For $x\geq 0$ the function
@@
Applying the last theorem, we get the following:

### v4

@@ Abstract.
$\frac{\pi\alpha_{1}}{2}<\arg\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}<\frac{\pi\alpha_{2}}{2},\quad(z\in\mathbb{D}).$
@@ 1. Introduction
On the other hand, strongly starlike functions are a subset of starlike functions, where the condition is stronger, ensuring a more restricted behavior of the image of the unit disk under the function. These functions are of interest in complex analysis and have applications in various areas, including the study of univalent functions and geometric function theory.
@@ 2. Preliminaries
Consider the set $\mathcal{H}$ consisting of functions $f$ that are holomorphic in the open unit disk $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$. Furthermore, let $\mathcal{A}$ be a subset of $\mathcal{H}$ consisting of functions $f$ that can be represented in the form:
@@
These functions in $\mathcal{A}$ are subject to a normalization condition, namely, $f(0)=0$ and $f^{\prime}(0)=1$, when evaluated within the unit disk $\mathbb{D}$. The subclass of $\mathcal{A}$ consisting of all univalent functions $f$ in $\mathbb{D}$ will be denoted by $\mathcal{S}$. We say that a function $f\in\mathcal{S}$ is starlike if and only if
@@
In the open unit disk $\mathbb{D}$, two analytic functions, denoted as $f$ and $g$, are considered in a subordinate relationship if $f(z)\prec g(z)$ or simply $f\prec g$. This relationship holds when there exists a Schwarz function $w$ defined in $\mathbb{D}$, satisfying the initial conditions $w(0)=0$ and $|w(z)|\leq 1$, and such that $f(z)=g(w(z))$ for all $z\in\mathbb{D}$.
In this paper, we consider the following analytic function
(2.2) $G(z):=G(\alpha_{1},\alpha_{2},c)(z)=\left(\frac{1+cz}{1-z}\right)^{(\alpha_{1}+\alpha_{2})/2},\quad(z\in\mathbb{D}),$
@@
$\left\{w\in\mathbb{C}:-\frac{\pi\alpha_{1}}{2}<\arg\,w<\frac{\pi\alpha_{2}}{2}\right\},$
see [35, Lemma 2] for more details. It is also worth mentioning that this function is a special case of the function defined in [1, (2.2)], where $x=c$ and
(2.3) $\alpha=\frac{\alpha_{1}+\alpha_{2}}{2}.$
It follows from [1, (2.2)] that
$G(z)=1+\sum_{n=1}^{\infty}A_{n}z^{n},$
@@
(2.4) $A_{n}:=A_{n}(\alpha_{1},\alpha_{2},c)=\sum_{k=1}^{n}\binom{n-1}{k-1}\binom{(\alpha_{1}+\alpha_{2})/2}{k}(1+c)^{k},\quad(n\geq 1).$
It is easy to see that
[…]
where $\alpha$ is given by (2.3). Moreover, by [1, (2.3)] we have
$A_{n}(\alpha_{1},\alpha_{2},c)=\frac{\alpha_{1}+\alpha_{2}}{2}(1+c)\,B_{n}(\alpha_{1},\alpha_{2},c),\quad(n\geq 1),$
where
(2.6) $B_{n}(\alpha_{1},\alpha_{2},c):={}_{2}F_{1}\!\left(1-n,\,1-\frac{\alpha_{1}+\alpha_{2}}{2};\,2;\,1+c\right),\quad(n\geq 1).$
@@
$_{2}F_{1}(a,b;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{(c)_{n}(1)_{n}}z^{n},\quad(a,b,c\in\mathbb{C},c\not\in\{0,-1,-2,\ldots\},z\in\mathbb{D}).$
The expression $(a)_{n}$, denoted as the Pochhammer symbol, signifies the product of consecutive terms starting from $a$ and ending at a+n1a+n-1, i.e., $(a)_{n}=a(a+1)\cdots(a+n-1)$ with the convention that $(a)_{0}$ is equal to $1$. For a deeper exploration of $G$ and its properties, refer to the comprehensive discussion in reference [1, Section 2].
@@
$|A_{n}|\leq\lambda\,|B_{n}(\alpha_{1},\alpha_{2},1)|,$
where
(2.7) $\lambda=(\alpha_{1}+\alpha_{2})\cos\left(\frac{\pi\theta}{2}\right)\quad\text{with}\quad\theta=\frac{\alpha_{2}-\alpha_{1}}{\alpha_{2}+\alpha_{1}}.$
The main purpose of this paper is to study the class $\mathcal{S}^{*}(\alpha_{1},\alpha,_{2})$, which is provided below:
@@
(2.8) $\frac{\pi\alpha_{1}}{2}<\arg\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}<\frac{\pi\alpha_{2}}{2},\quad(z\in\mathbb{D}),$
@@
The introduction of the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ is attributed to Takahashi and Nunokawa, specifically in the context of inequalities (2.8), see [33]. We recall here the fact that in [7] and in [8] a similar class was studied. It is clear that $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})\subset\mathcal{S}^{*}$ and that $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ is a subclass of the class of strongly starlike functions of order $\beta=\max\{\alpha_{1},\alpha_{2}\}$, i.e. $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})\subset\mathcal{S}^{*}(\beta,\beta)\equiv\mathcal{SS}^{*}(\beta)$.
@@
In the following, we recall some other useful lemmas.
Lemma 2.4 (See [25]).
@@
$|w_{2}-tw_{1}^{2}|\leq\left\{\begin{array}[]{ll}-t,&\hbox{$t\leq-1$;}\\ \\ 1,&\hbox{$-1\leq t\leq 1$;}\\ \\ t,&\hbox{$t\geq 1$.}\end{array}\right$.
which simplifies to $|w_{2}+tw_{1}^{2}|\leq\max\{1,|t|\}$ for $t\in\mathbb{R}$. All inequalities are sharp.
The following lemma is due to Prokhorov and Szynal.
@@
(See [19]) If $w(z)=\sum_{n=1}^{\infty}w_{n}z^{n}$ is a Schwarz function, then for any complex numbers $\rho$ and $\tau$ the following sharp estimate holds:
@@
[…]
@@ 3. Main Results
In this paper, we find a general upper bound for $|a_{n}|$ of the function $f\in\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ and $f\in\mathcal{SS}^{*}(\beta)$, where $n=2,3,\ldots$. Our findings generally do not produce sharp results, except when $n=2$. Of course, in Theorem 3.4 below, we obtain sharp estimates for the first three coefficients of $f\in\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$.
@@
Let $f$ be of the form (2.1) belonging to the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$, where $0<\alpha_{1},\alpha_{2}\leq 1$. Then $|a_{2}|\leq\lambda$ and for $n\geq 3$,
(3.1) $|a_{n}|\leq\frac{\lambda}{n-1}\prod_{k=2}^{n-1}\left(1+\frac{\lambda}{k-1}\right),$
where $\lambda$ is given by (2.7). The bound is sharp for $n=2$.
@@
(3.4) $|Q_{n}|\leq|A_{1}|=(\alpha_{1}+\alpha_{2})\cos\left(\frac{\pi\theta}{2}\right)=\lambda,\quad(n=1,2,3,\ldots).$
@@
$\leq\frac{\lambda}{n-1}(|a_{1}|+|a_{2}|+\cdots+|a_{n-1}|)$
$=\frac{\lambda}{n-1}\sum_{k=1}^{n-1}|a_{k}|.$
It is clear that $|a_{2}|\leq\lambda$. To prove the remaining part of the theorem, we need to show that
(3.5) $\frac{\lambda}{n-1}\sum_{k=1}^{n-1}|a_{k}|\leq\frac{\lambda}{n-1}\prod_{k=2}^{n-1}\left(1+\frac{\lambda}{k-1}\right),$
for $n=3,4,\ldots$. Using induction and simple calculation, we could prove the inequality (3.5). Hence, the desired estimate for $|a_{n}|$, $(n=3,4,5,\ldots)$, follows as asserted in (3.1).
It is easy to see that the extremal function
(3.6) $f(z)=z\exp\!\left(\int_{0}^{z}\frac{G(w(\zeta))-1}{\zeta}\,d\zeta\right),$
gives $|a_{2}|=\lambda$ exactly, where $w$ is a Schwarz function. For $n\geq 3$, the bounds are generally not sharp because of the compounding effect of the subordination chain. This completes the proof of the theorem. ∎
@@
where $0<\beta\leq 1$. The bound is sharp for $n=2$.
@@
Consider a function $f\in\mathcal{A}$, which takes the form (2.1) and belongs to the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$, where $\alpha_{1},\alpha_{2}\in(0,1]$. Let $\lambda$ be defined by (2.7), and $\gamma=(3\alpha-1)/2$, where $\alpha$ is given by (2.3). Then
$|a_{2}|\leq\lambda,\qquad|a_{3}|\leq\left\{\begin{array}[]{ll}\frac{\lambda}{2},&\qquad\hbox{$0<\alpha\leq 1/3$;}\\ \\ \frac{\lambda}{2}\sqrt{1+4\gamma(1+\gamma)\cos^{2}\!\left(\frac{\pi\theta}{2}\right)},&\qquad\hbox{$1/3\leq\alpha\leq 1$,}\end{array}\right$.
@@
$|a_{4}|\leq\frac{\lambda}{3}\,H(\rho,\tau),$
where $H(\rho,\tau)$ is given by (2.11) with
(3.7) $\rho=\frac{3}{2}A_{1}+\frac{2A_{2}}{A_{1}},\quad\tau=\frac{1}{2}A_{1}^{2}+\frac{3}{2}A_{2}+\frac{A_{3}}{A_{1}}.$
Each inequality is sharp.
@@
where $A_{n}$, $n=1,2,3$, are defined in (2.5). Equating the coefficients of the corresponding terms in the last relation leads to
@@
It follows from the first equality in (3.9) that $|a_{2}|=|A_{1}||w_{1}|\leq|A_{1}|=\lambda$, which yields the sharp upper bound for $|a_{2}|$.
The second coefficient relation in (3.9) can be rewritten as
$2a_{3}=A_{1}\left(w_{2}+\left(\frac{A_{2}}{A_{1}}+A_{1}\right)w_{1}^{2}\right).$
An application of Lemma 2.5 yields
(3.11) $2|a_{3}|\leq|A_{1}|\max\left\{1,\left|\frac{A_{2}}{A_{1}}+A_{1}\right|\right\}.$
A direct computation using (2.5) shows that
$\frac{A_{2}}{A_{1}}+A_{1}=1+\frac{3\alpha-1}{2}(1+c)=1+\gamma(1+c).$
Taking moduli gives
$\left|\frac{A_{2}}{A_{1}}+A_{1}\right|=\sqrt{1+4\gamma(1+\gamma)\cos^{2}\!\left(\frac{\pi\theta}{2}\right)}.$
It is easy to see that $|A_{2}/A_{1}+A_{1}|\geq 1$ if and only if $\gamma\geq 0$, which implies that $\alpha\geq 1/3$. Therefore, we have
$\max\left\{1,\left|\frac{A_{2}}{A_{1}}+A_{1}\right|\right\}=\sqrt{1+4\gamma(1+\gamma)\cos^{2}\!\left(\frac{\pi\theta}{2}\right)}.$
On the other hand, $|A_{2}/A_{1}+A_{1}|\leq 1$ if and only if $\gamma\leq 0$, which implies that $\alpha\leq 1/3$. Thus,
$\max\left\{1,\left|\frac{A_{2}}{A_{1}}+A_{1}\right|\right\}=1,$
if $0<\alpha\leq 1/3$. Consequently, by (3.11) we obtain the sharp bound for $|a_{3}|$.
From the coefficient relation (3.10)
@@
$\rho=\frac{3}{2}A_{1}+\frac{2A_{2}}{A_{1}},\quad\tau=\frac{1}{2}A_{1}^{2}+\frac{3}{2}A_{2}+\frac{A_{3}}{A_{1}}.$
@@
$|a_{4}|\leq\frac{|A_{1}|}{3}\,H(\rho,\tau)=\frac{\lambda}{3}\,H(\rho,\tau),$
where $H(\rho,\tau)$ is given by (2.11).
The coefficient inequalities in Theorem 3.4 are attained by different extremal functions. More precisely, each bound is realized by choosing an appropriate extremal Schwarz function $w$ in the subordination relation
(3.12) $\frac{zf^{\prime}(z)}{f(z)}=G(w(z)),$
where $G$ is given by (2.2). For any such Schwarz function $w$, the corresponding extremal function is uniquely determined by the extremal function (3.6). The sharp bound for $|a_{2}|$ is attained when $w(z)=e^{i\phi}z$ for some $\phi\in\mathbb{R}$. Therefore, the corresponding extremal function is given by:
$f_{2}(z)=z+A_{1}e^{i\phi}z^{2}+\frac{A_{2}+A_{1}^{2}}{2}e^{2i\phi}z^{3}+\frac{A_{3}+2A_{1}A_{2}+A_{1}^{3}}{3}e^{3i\phi}z^{4}+\cdots.$
The sharpness of the bound for $|a_{3}|$ depends on the range of the parameter $\alpha$. If $0<\alpha\leq 1/3$, then
$\max\left\{1,\left|\frac{A_{2}}{A_{1}}+A_{1}\right|\right\}=1,$
and equality in Lemma 2.5 is attained for Schwarz functions of the form $w(z)=e^{i\psi}z^{2}$, $\psi\in\mathbb{R}$. The corresponding extremal function satisfies
$f_{31}(z)=z+\frac{A_{1}}{2}e^{i\psi}z^{3}+O(z^{5}).$
If $1/3\leq\alpha\leq 1$, then
$\max\left\{1,\left|\frac{A_{2}}{A_{1}}+A_{1}\right|\right\}=\left|\frac{A_{2}}{A_{1}}+A_{1}\right|,$
and equality is attained for Schwarz functions of the form $w(z)=e^{i\phi}z$, $\phi\in\mathbb{R}$. In this case, the extremal function is given by
$f_{32}(z)=z+A_{1}e^{i\phi}z^{2}+\frac{A_{2}+A_{1}^{2}}{2}e^{2i\phi}z^{3}+\cdots.$
Finally, the sharp bound for $|a_{4}|$ is attained by Schwarz functions realizing equality in the Prokhorov–Szynal lemma, which are finite Blaschke products of degree at most two. In particular, equality is attained for Schwarz functions of the form
$w(z)=e^{i\phi}\frac{z-a}{1-\overline{a}z},\quad|a|\leq 1,$
for which equality holds in the Prokhorov–Szynal lemma.
Consequently, there is no single extremal function that simultaneously attains equality in all three coefficient bounds. This completes the proof. ∎
@@
We illustrate the bound for $|a_{4}|$ in Theorem 3.4 for several special choices of the parameters $\alpha_{1}$ and $\alpha_{2}$.
Case 1. Let $\alpha_{1}=\alpha_{2}=1$. Then $\theta=0$, $c=1$, and $\alpha=(\alpha_{1}+\alpha_{2})/2=1$. Using (2.5), we obtain
$A_{1}=2,\qquad A_{2}=2,\qquad A_{3}=2.$
Consequently,
$\rho=\frac{3}{2}A_{1}+\frac{2A_{2}}{A_{1}}=5,\qquad\tau=\frac{1}{2}A_{1}^{2}+\frac{3}{2}A_{2}+\frac{A_{3}}{A_{1}}=6.$
Since $(\rho,\tau)\in\Omega_{7}$, Lemma 2.6 yields $H(\rho,\tau)=|\rho|=5$. Therefore,
$|a_{4}|\leq\frac{\lambda}{3}H(\rho,\tau)=\frac{2}{3}\cdot 5=\frac{10}{3}.$
Case 2. Let $\alpha_{1}=\alpha_{2}=\tfrac{1}{2}$. Then $\theta=0$, $c=1$, and $\alpha=\tfrac{1}{2}$. From (2.5), we find
$A_{1}=1,\qquad A_{2}=\tfrac{1}{2},\qquad A_{3}=\tfrac{1}{2}.$
Hence,
$\rho=\frac{3}{2}A_{1}+\frac{2A_{2}}{A_{1}}=\frac{5}{2},\qquad\tau=\frac{1}{2}A_{1}^{2}+\frac{3}{2}A_{2}+\frac{A_{3}}{A_{1}}=\frac{7}{4}.$
Since $(\rho,\tau)\in\Omega_{6}$, Lemma 2.6 gives $H(\rho,\tau)=|\rho|=\tfrac{5}{2}$. Thus,
$|a_{4}|\leq\frac{\lambda}{3}H(\rho,\tau)=\frac{1}{3}\cdot\frac{5}{2}=\frac{5}{6}.$
Case 3. Let $\alpha_{1}=1$ and $\alpha_{2}=\tfrac{1}{2}$. Then $\alpha=\tfrac{3}{4}$ and $\theta=-\tfrac{1}{3}$, so that $c=e^{-i\pi/3}$ and $|1+c|=\sqrt{3}$. Using (2.5), we obtain
$A_{1}=\tfrac{3}{4}(1+c),\quad A_{2}=\tfrac{3}{4}(1+c)-\tfrac{3}{32}(1+c)^{2},\quad A_{3}=\tfrac{3}{4}(1+c)-\tfrac{3}{16}(1+c)^{2}+\tfrac{5}{128}(1+c)^{3}.$
A direct computation yields
$\rho=\frac{3}{2}A_{1}+\frac{2A_{2}}{A_{1}}\approx 3.79,\qquad\tau=\frac{1}{2}A_{1}^{2}+\frac{3}{2}A_{2}+\frac{A_{3}}{A_{1}}\approx 3.02.$
Since $(\rho,\tau)\in\Omega_{6}$, Lemma 2.6 again implies $H(\rho,\tau)=|\rho|$. Hence,
$|a_{4}|\leq\frac{\lambda}{3}H(\rho,\tau)\approx 1.64.$
Remark on boundary behavior. If $\alpha_{1}+\alpha_{2}\to 0$, then $A_{1}\to 0$, and the parameters $\rho$ and $\tau$ are no longer well defined. Therefore, such limiting cases fall outside the scope of Theorem 3.4 and are excluded from consideration.
In all admissible cases, the bound for $|a_{4}|$ is sharp in the sense of Lemma 2.6, and equality is attained by extremal Schwarz functions given by finite Blaschke products of degree at most two.
3.1. Logarithmic Coefficients.
@@
$|\gamma_{1}|\leq 1,\quad{\rm and}\quad|\gamma_{2}|\leq\frac{1}{2}(1+2e^{-2})\approx 0.635\ldots,$
hold. For $n\geq 3$, estimating $\gamma_{n}$ is much more difficult, and no significant upper bounds for $|\gamma_{n}|$ are known when $f\in\mathcal{S}$; the problem remains open for $n\geq 3$. The sharp upper bounds for the modulus of logarithmic coefficients have been established for only a limited number of function subclasses within $\mathcal{S}$; the class $\mathcal{S}^{*}$ offers a relatively straightforward proof that $|\gamma_{n}|\leq 1/n$ for $n\geq 1$, with equality observed in the Koebe function. For other subclasses within $\mathcal{S}$, further details can be found in references [10, 20, 21, 34].
In what follows, we focus on estimating the logarithmic coefficients of functions belonging to the class $\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$. First, we need the following theorem:
@@
where G(z)1G(z)-1 is convex univalent. For $x\geq 0$ the function
@@
Applying the last theorem, we get the following.